OpenStudy (anonymous):

cosx (cos(-x) - tan(-x)) = cosx (cos(x) + tan(x) Why does cos(x) become positive?

OpenStudy (anonymous):

Because cosine is an even function.

OpenStudy (anonymous):

A function f(x) is even if f(-x) = f(x).

OpenStudy (anonymous):

In the case of tangent, you have \[+\tan (-x)\\ +\frac{\sin (-x)}{\cos (-x)}\\ +\frac{-\sin x}{\cos x}\;\;\;\;\;\;\;\;\text{(because sine is an odd function)}\\ -\frac{\sin x}{\cos x}\\ -\tan x\] And for odd functions: A function f(x) is odd if f(-x) = -f(x).

OpenStudy (anonymous):

oh yes i forgot, thanks